OpenStudy (anonymous):

find the domain and sketch the graph of the function G(x)=3x+IxI/x

OpenStudy (anonymous):

Do you know how to calculate the domain of that function?that is how your graph looks like apprxomately.

OpenStudy (anonymous):

plz explain

OpenStudy (anonymous):

The domain is every possible value x can have. In your example, x can be any number except for 0 because you cannot divide by zero |x|/x So the domain of your function is every Real number other than zero, this can be written as \[\left\{ x \epsilon R | x \neq 0 \right\}\]

OpenStudy (anonymous):

or\[\left\{ - \infty < x < 0 and 0 < x < \infty \right\}\] Depending on how your teacher wants it

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

yeas thnks